Two countries produce and consume T-shirts: the US and the ROW. Problems 1-2 are based on the supply and demand schedules for the two countries given below. Note: The supply and demand curves are straight lines. Quantities are in millions of T-shirts. US ROW 32 13 28 26 10 20 18 12 12 13 14 15
This problem asks you to examine the welfare effects of opening trade between the two countries. Please draw new graphs (separate from question I). a. Label the area in your graph that represents consumer surplus in the ROW before trade. What is b. Label the area in your graph that represents producer surplus in the ROW before trade. What is the S value of consumer surplus? the $ value of producer surplus? Label the area in your graph that represents consumer surplus in the ROW after the opening of trade with the US. What is the S value of consumer surplus? Label the area in your graph that represents producer surplus in the ROW after the opening of trade with the US. What is the S value of producer surplus? How much does the ROW gain from trade? ROW consumers? ROW producers? Give S values. d.
Answer #1

Part a

Consumer surplus is the difference between what consumers are willing to pay and consumers are actually paying, i.e., equilibrium price. In this case it is 6 and the consumer surplus is given by the shaded area in the below image.

Dollar value of consumer surplus = area of the shaded area (triangle)

= 0.5 X base of triangle X height of triangle

= 0.5 X 7 X (13-6) = 24.5 million $

Consumer surplus of ROW before trade 16 14 12 6 14 16 18 10 12

Part b

Producer surplus is the difference between what sellers are actually receiving, i.e., equilibrium price and what sellers are willing to receive. In this case it is 6 and the producer surplus is given by the shaded area in the below image.

Dollar value of producer surplus = area of the shaded area (triangle + rectangle)

= (0.5 X base of triangle X height of triangle) + (length X breadth)

= (0.5 X (7-1) X 6) + (6 X 1) = 18 + 6 million $ = 24 million $

Producer surplus of ROW before trade 16 14 12 10 6 10 12 14 16 18

Part c

Using the demand and supply schedules for ROW and US, we arrive at the world demand and supply schedule at each price level using horizontal summation. For example, at price level 8, world quantity supply = 4 + 7 = 11 and world quantity demand = 16 + 5 =21. Below is the world demand and supply schedule and the the shaded area is the consumer surplus

Dollar value of consumer surplus = area of the shaded area (triangle)

= 0.5 X base of triangle X height of triangle

= 0.5 X 10 X (45-15) = 150 $

P (world) S (world) D (world) 45 42 39 36 0 Consumer surplus of world after trade 2 45 2 4 35 30 30 27 25 24 20 21 15 1810 15

Part D

Using the same schedule in part C, the shaded area in the below image would be the world producer surplus.

Dollar value of producer surplus = area of the shaded area (triangle + rectangle)

= (0.5 X base of triangle X height of triangle) + (length X breadth)

= (0.5 X (10-1) X 15) + (15 X 1) = 67.5 + 15 million $ = 82.5 million $

P (world) S (world) D (world 45 42 39 36 0 Producer surplus of world after trade 0 50 45 2 2 4 35 30 30 27 25 24 20 21 15 18